# Retrieval-Augmented Generation; Entity Linking; Domain Adaptation; Wikidata; Educational Question Answering

### v1

@@ 1 Introduction
In the current technological landscape, dominated by Large Language Models (LLMs), the way we approach information and data has profoundly changed. These models are capable of producing fluent and context-aware text across a wide range of domains, making them powerful tools for knowledge access, task automation, and decision support. Their impressive generative capabilities, however, come with inherent limitations. Despite these strengths, LLMs are not always the best way to access knowledge, especially in critical, specialized or high-precision domains, because they can sometimes provide incorrect, inconsistent or unverifiable information. This phenomenon - commonly referred to as hallucination Huang2025 - poses significant risks when accuracy and reliability are essential, such as in scientific research  Zhang_2025, medicine Asgari2025; Pal_2023, and education Qian2026; Vrdoljak2025.
@@ 1.1 Related Work
To improve the factual grounding of large language models (LLMs), several studies have examined Retrieval-Augmented Generation (RAG) systems as a promising framework to address this open research challenge. The original architecture proposed by Lewis et al. (2020) (lewis2020retrieval) integrates a neural retriever with a generative model, enabling language models to access external knowledge sources dynamically. A standard RAG pipeline can be conceptually divided into two main components: a retrieval phase and a generation phase. The retrieval component is often based on semantic similarity between query and document passages, although alternative approaches have been explored. For example, Mongiovì et al. (2024) (mongiovi2024graal) introduced GRAAL, a graph-based retrieval system that collects related passages across multiple documents by exploiting inter-entity relationships. A wide range of retrieval strategies have been proposed to improve access to relevant information; however, many still face challenges in capturing fine-grained semantic meaning, particularly in specialized domains that contain ambiguous or domain-specific terminology.
@@
Entity Linking (EL) has been proposed as a method to improve the disambiguation of named entities by aligning textual mentions with entries in structured knowledge bases such as Wikidata (wikidata2025) or DBpedia. Shen et al. (2015) (shen2014entity) provided an extensive overview of EL challenges and techniques. Möller et al. (2022) (moller2022survey) presented a comprehensive survey of Wikidata-based EL methods and datasets. Several frameworks have been developed in this area, including BLINK (wu2019scalable), ReLiK (orlando2024relik), and OpenTapioca (delpeuch2019opentapioca), each adopting different architectures and disambiguation strategies. Most current EL systems, however, have been primarily developed for the English language, which limits their direct applicability in multilingual or non-English domains.
@@
At the same time, Educational Artificial Intelligence (EAI) has become a rapidly expanding field of research. AI-based systems are increasingly applied to support learning, personalize content delivery, and facilitate knowledge acquisition. Mageira et al. (2022) (mageira2022educational) discussed the pedagogical potential of educational chatbots designed for content and language-integrated learning. More recently, Swacha et al.(2025) (swacha2025retrieval) and Li et al. (2025) (li2025retrieval) provided a systematic survey of RAG applications in education, illustrating how retrieval-augmented approaches can enhance information access and automated question answering in academic contexts.
@@ 1.2 Research Objectives and Contributions
    We evaluate the efficacy of our proposed RRF-Based Re-ranking strategy by comparing it against a baseline Weighted-Score Re-ranking and a high-complexity RRF + Cross-Encoder Re-ranking.
@@
This paper is organized as follows: Section 2 details the methodology, describing the hybrid RAG architecture, the Entity Linking module, and the implemented re-ranking strategies. Section 3 presents the experimental setup, including the construction of the custom educational dataset, the benchmarks, and the evaluation metrics. Section 4 reports the quantitative and qualitative results obtained from the experiments. Finally, Section 5 discusses the findings and draws the main conclusions, highlighting the domain mismatch phenomenon and outlining limitations and future directions.
@@ 2.1 RAG baseline
Each chunk in the corpus was encoded into a fixed-length 1024-dimensional vector representation using the multilingual-e5-large model by Sentence Transformers  [^1] https://sbert.net, chosen for its strong multilingual retrieval capabilities and competitive performance on cross-lingual benchmarks. The resulting embeddings were normalized and stored in a FAISS index employing an inner-product 222for normalized vectors it is equivalent to cosine similarity similarity metric to enable efficient dense retrieval at query time. During inference, a user query is first encoded with the same embedding model, then the top-$K$ most similar vectors are probed and the corresponding chunks are retrived. These retrieved passages are concatenated to form a context window, which is then provided as additional input to the generative model.
@@ 2.2 Entity Linking Module
Every chunk in the dataset was pre-processed to extract named entities using the SpaCy pipeline (spacy2023). Specifically, we employed the it_core_news_lg  [^3] https://spacy.io/models/it#it_core_news_lg model, a large pre-trained pipeline for the Italian language trained on a massive corpus of news and media text, chosen for its superior accuracy in Named Entity Recognition (NER) compared to smaller variants.
@@
where $\alpha$ (=0.9=0.9) controls the balance between semantic and popularity signals, a value empirically chosen after a preliminary tuning phase. The candidate with the highest score is selected as the final linked entity and stored, together with all metadata and supporting information, in a JSON structure with the addition of linked entities.
@@ 2.3 Ranking Strategies and Integration into the Retrieval Pipeline
where $K=60$ is the standard smoothing constant. This approach was selected as the primary method for ELERAG because it robustly balances semantic relevance with factual entity matching without requiring the manual tuning of weights or the high computational cost of cross-encoders. As demonstrated in Section 4, this strategy yielded the best performance on our specialized educational dataset.
@@
To rigorously evaluate the effectiveness of our proposed strategy, we implemented two alternative re-ranking methods for comparison:
@@
  2. 2.
    RRF + Cross-Encoder Re-ranking. To benchmark our approach against a pure semantic SOTA method, we implemented a three-stage pipeline. After dense retrieval and RRF fusion, the top candidates are re-scored by a transformer-based Cross-Encoder (crossencoder) (mmarco-mMiniLMv2-L12-H384-v1). Unlike bi-encoders (like E5), the Cross-Encoder processes the query and document simultaneously, capturing finer semantic nuances at a higher computational cost. For this specific configuration, to maximize initial recall before the expensive scoring step, we retrieved a larger pool of candidates (Top-50) from the vector index and selected the Top-20 final chunks after re-ranking.
@@
Figure 1 illustrates the complete workflow of our proposed ELERAG method. The architecture processes the query in parallel streams—extracting entity-based features and computing dense embeddings—before fusing the results via the RRF module to feed the LLM. The other experimental configurations can be understood as variations of this schema: the Standard RAG baseline utilizes only the lower branch (Dense Embedding $\rightarrow$ Vector Index $\rightarrow$ LLM), bypassing the Entity Linking and Re-ranking stages. The Weighted-Score Re-ranking configuration replaces the RRF block with the linear combination strategy described above. Finally, the RRF + Cross-Encoder Re-ranking configuration adds a further refinement block between the RRF stage and the LLM generation.
@@ 3.1 Educational Data
The dataset used in this study was constructed from two Italian university courses, namely Applied Economics and Language and Communication, offered by an Italian telematic university. Each course consisted of several video lectures, totalling 50 classes over 32 hours of material, which were first transcribed into text using the Whisper Turbo automatic speech recognition model (radford2022whisper). The transcription process also generated precise word-level timestamps, which were later used to associate each text segment with its corresponding temporal interval in the original lecture.
To prepare the textual material for retrieval, the transcribed lectures were segmented into coherent chunks using the SpaCy Python library (spacy2023). To ensure semantic coherence, segmentation was strictly constrained to sentence boundaries: sentences were grouped to form chunks with a length between 20 and 300 tokens. This strategy prevents cutting sentences in the middle, preserving the syntactic and semantic integrity of the text. For each chunk, start and end timestamps were retained. All chunks were stored in JSON format and enriched with named entities automatically extracted by SpaCy, which were later used by the EL module described in Section 2.2.
@@ 3.2 Evaluation setup
After generation, the dataset underwent a two-step validation process. First, a secondary GPT-4o prompt was used to verify the correctness and consistency between questions and their corresponding answers. Subsequently, a manual validation was performed by human annotators to identify and filter out ambiguous, ill-formed, or hallucinated questions that might have bypassed the automated check. This rigorous process ensured the high quality of the final question set used for evaluation.
@@ 3.3 Evaluation Metrics
  • Subjective metrics: Completeness, relevance, and clarity, each evaluated by gpt-4o on a 1–10 scale for the generated answers obtained by concatenating the top-3 retrieved chunks. Although LLM-based evaluation cannot fully replace human judgment, it provides a scalable and replicable proxy for consistently estimating answer quality across the benchmark.
@@ 3.4 Evaluation methods
    Method 2: Subjective evaluation through LLM-based scoring. To assess the quality of the generated text, final answers were produced by the RAG system using the top-$K$ retrieved chunks as context. For this specific evaluation, $K$ was set to 3 across all configurations to ensure comparable context lengths. Subsequently, a separate gpt-4o instance was prompted to act as an external evaluator, scoring the generated answers on a 1–10 scale according to three criteria: completeness, relevance, and clarity.
@@ 4.1 Method 1: Classical Evaluation on Retrieval
Analysis of Results.
The experimental data confirms that integrating Entity Linking significantly improves retrieval precision. All entity-enhanced pipelines outperformed the baseline in terms of Exact Match (EM), indicating a superior ability to place the precise gold chunk at the very top of the ranking.
@@
Comparison of Ranking Strategies. Our proposed ELERAG strategy emerged as the most effective method for this domain, achieving the highest scores in EM ($0.565$), Precision@1 ($0.696$), and MRR_gold ($0.668$) and MRR_rel_docs ($0.779$). This superiority demonstrates that fusing the entity signal with dense retrieval via RRF is more robust than a simple linear combination (Weighted-Score), effectively filtering out semantically similar but factually incorrect chunks.
Cross-Encoder vs. Entity Signal. Comparing our method with the RRF + Cross-Encoder reveals a critical insight consistent with the Domain Mismatch hypothesis. The Cross-Encoder achieved the highest Recall@1 ($0.408$), proving its strong capability to identify relevant content in a broad sense. However, it failed to translate this into superior Exact Match or MRR scores, performing worse than ELERAG on these precision metrics. This suggests that while the generic, pre-trained Cross-Encoder captures general semantics well, it lacks the specific domain grounding provided by the explicit Entity Linking signal. Consequently, ELERAG proves to be not only more accurate for the top-ranking position but also significantly more computationally efficient than the Cross-Encoder pipeline.
Recall Trade-off. It is worth noting that the Baseline maintains higher Recall@5 and Recall@10 scores compared to the re-ranking methods. This is a typical behavior of re-ranking architectures, which aggressively optimize the top positions (Top-1 to Top-3) to maximize precision for the limited context window of the LLM. While this may push some marginally relevant documents further down the list (slightly affecting R@10), the substantial gain in MRR and EM is far more valuable for a RAG system, where the generator’s performance depends primarily on the quality of the very first retrieved chunks.

### v2

@@ 1 Introduction
In the current technological landscape, dominated by Large Language Models (LLMs), the way we approach information and data has profoundly changed. These models are capable of producing fluent and context-aware text across a wide range of domains, making them powerful tools for knowledge access, task automation, and decision support. Their impressive generative capabilities, however, come with inherent limitations. Despite these strengths, LLMs are not always the best way to access knowledge, especially in critical, specialized or high-precision domains, because they can sometimes provide incorrect, inconsistent or unverifiable information. This phenomenon - commonly referred to as hallucination Huang2025 - poses significant risks when accuracy and reliability are essential, such as in scientific research  Zhang_2025, medicine Asgari2025; Pal_2023; haq2026heal, and education Qian2026; Vrdoljak2025.
@@ 1.1 Related Work
To improve the factual grounding of large language models (LLMs), several studies have examined Retrieval-Augmented Generation (RAG) systems as a promising framework to address this open research challenge. The original architecture proposed by Lewis et al. (2020) (lewis2020retrieval) integrates a neural retriever with a generative model, enabling language models to access external knowledge sources dynamically. A standard RAG pipeline can be conceptually divided into two main components: a retrieval phase and a generation phase.
Historically, the retrieval component has relied on sparse, lexical matching techniques, with BM25 (robertson2009probabilistic) standing as the most widely adopted probabilistic model based on exact keyword overlap. While highly effective for specific terminology, sparse methods struggle with the vocabulary mismatch problem. Consequently, recent RAG architectures have heavily shifted towards dense retrieval, utilizing neural embeddings based on semantic similarity to capture deep relationships. Alternative retrieval approaches have also been explored. For example, Mongiovì et al. (2024) (mongiovi2024graal) introduced GRAAL, a graph-based retrieval system that collects related passages across multiple documents by exploiting inter-entity relationships.
Despite these advancements, many dense retrieval models still face challenges in capturing fine-grained semantic meaning, particularly in specialized domains that contain ambiguous or domain-specific terminology Ji2023. Because dense models often underperform on exact string matching for rare entities or acronyms, state-of-the-art approaches increasingly employ hybrid search strategies weinberg2026hybrid. These mixed methods combine the exact lexical precision of sparse models (like BM25) with the semantic generalization of dense embeddings, typically merging the distinct retrieval lists through aggregation algorithms such as Reciprocal Rank Fusion (RRF) (rrf).
@@
Our work extends this hybrid retrieval paradigm by replacing the purely lexical sparse retriever with an explicit, knowledge-grounded signal. Specifically, Entity Linking (EL) has been proposed as a method to improve the disambiguation of named entities by aligning textual mentions with entries in structured knowledge bases such as Wikidata (wikidata2025) or DBpedia. Shen et al. (2015) (shen2014entity) provided an extensive overview of EL challenges and techniques. Möller et al. (2022) (moller2022survey) presented a comprehensive survey of Wikidata-based EL methods and datasets. Several frameworks have been developed in this area, including BLINK (wu2019scalable), ReLiK (orlando2024relik), and OpenTapioca (delpeuch2019opentapioca), each adopting different architectures and disambiguation strategies. Most current EL systems, however, have been primarily developed for the English language, which limits their direct applicability in multilingual or non-English domains.
@@
At the same time, Educational Artificial Intelligence (EAI) has become a rapidly expanding field of research. AI-based systems are increasingly applied to support learning, personalize content delivery, and facilitate knowledge acquisition Furini202298. Mageira et al. (2022) (mageira2022educational) discussed the pedagogical potential of educational chatbots designed for content and language-integrated learning. More recently, Swacha et al.(2025) (swacha2025retrieval) and Li et al. (2025) (li2025retrieval) provided a systematic survey of RAG applications in education, illustrating how retrieval-augmented approaches can enhance information access and automated question answering in academic contexts. More broadly, recent studies have explored the integration of LLM-based architectures within domain-specific decision-support systems, including frameworks that combine language models with multimodal data processing and explainable reasoning for applications such as healthcare decision making HAQ2026102696.
@@ 1.2 Research Objectives and Contributions
    We evaluate the efficacy of our proposed RRF-Based Re-ranking strategy by comparing it against a baseline Weighted-Score Re-ranking, a high-complexity RRF + Cross-Encoder Re-ranking and a standalone Cross-Encoder Re-ranking.
@@
This paper is organized as follows: Section 2 details the methodology, describing the hybrid RAG architecture, the Entity Linking module, and the implemented re-ranking strategies. Section 3 presents the experimental setup, including the construction of the custom educational dataset, the benchmarks, and the evaluation metrics. Section 4 reports the quantitative and qualitative results obtained from the experiments. Finally, Section 4.5 discusses the findings and draws the main conclusions, highlighting the domain mismatch phenomenon and outlining limitations and future directions.
@@ 2.1 RAG baseline
Each chunk in the corpus was encoded into a fixed-length 1024-dimensional vector representation using the multilingual-e5-large model by Sentence Transformers  [^1] https://sbert.net, chosen for its strong multilingual retrieval capabilities and competitive performance on cross-lingual benchmarks. The resulting embeddings were normalized and stored in a FAISS index employing an inner-product 222for normalized vectors it is equivalent to cosine similarity similarity metric to enable efficient dense retrieval at query time. During inference, a user query is first encoded with the same embedding model as the corpus; then, the top-$K$ most similar vectors are retrieved, and the corresponding document chunks are returned. These retrieved passages are concatenated to form a context window, which is then provided as additional input to the generative model.
@@ 2.2 Entity Linking Module
Every chunk in the dataset was pre-processed to extract named entities using the SpaCy pipeline (spacy2023). Specifically, we employed the it_core_news_lg  [^3] https://spacy.io/models/it#it_core_news_lg model, a large pre-trained pipeline for the Italian language trained on a massive corpus of news and media text, chosen for its superior accuracy in Named Entity Recognition (NER) compared to smaller variants. Since the corpus originates from automatically transcribed lecture speech, entity mentions may occasionally appear in abbreviated forms or lexical variants. The linking step partially mitigates this issue by relying on the Wikidata knowledge base, which maintains extensive lists of aliases and alternative labels for each entity. As a result, common variations (e.g., surnames, acronyms, or shortened forms) can still be resolved to the correct Wikidata identifier during candidate retrieval. In addition, the upstream transcription stage uses the Whisper Turbo ASR model, which provides high accuracy on proper names in lecture-style speech. This reduces the frequency of phonetic transcription errors that could otherwise affect the subsequent NER and linking stages.
@@
where $\alpha$ (=0.9=0.9) controls the balance between semantic and popularity signals. Note that the similarity score considers the whole mention context, therefore it is able to distinguish among multiple entities which correspond to the same mention.
The candidate with the highest score is selected as the final linked entity and stored, together with all metadata and supporting information, in a JSON structure with the addition of linked entities.
Consequently, if an entity mention is not detected due to transcription noise or spelling variations, the entity-based signal is simply absent. In such cases, the system naturally falls back to the dense semantic retrieval component, ensuring that the overall retrieval process remains robust.
@@ 2.3 Ranking Strategies and Integration into the Retrieval Pipeline
where K=60K=6044460 is the value suggested by the authors of RRF and is a de facto standard (rrf; tang2024found) is the standard smoothing constant. This approach was selected as the primary method for ELERAG because it robustly balances semantic relevance with factual entity matching without requiring the manual tuning of weights or the high computational cost of cross-encoders. As demonstrated in Section 4, this strategy yielded the best performance on our specialized educational dataset.
An intrinsic architectural advantage of this specific RRF implementation is its graceful degradation. In edge cases where the user query contains no recognizable named entities, the entity matching score evaluates to zero for all candidate chunks. Consequently, the entity-based sorting step preserves the initial order, meaning the entity rank ($r_{EL}$) simply mirrors the dense rank ($r_{dense}$). The final RRF calculation smoothly proceeds without altering the relative ordering, allowing the system to seamlessly fall back to the Standard Dense RAG baseline without computational interruptions or penalty assignments.
@@
To rigorously evaluate the effectiveness of our proposed strategy, we implemented three alternative re-ranking methods for comparison:
@@
    Hybrid Sparse-Dense Baseline. To thoroughly evaluate the contribution of the Entity Linking module against traditional exact-match approaches, we implemented a strong hybrid search baseline. This strategy combines the semantic representations of the multilingual-e5-large dense retriever with the exact keyword matching of a standard BM25 (robertson2009probabilistic) sparse retriever (via the rank_bm25 library). To ensure a fair comparison with ELERAG, the sparse and dense signals were fused using the identical Reciprocal Rank Fusion (RRF) mechanism employed in our proposed pipeline.
  2. 2.
@@
  3. 3.
    RRF + Cross-Encoder Re-ranking. This configuration explores a multi-stage refinement pipeline. After retrieving an initial pool of candidates via dense search and fusing them through RRF, the highest-ranking chunks are re-scored by a transformer-based Cross-Encoder (crossencoder) (namely, mmarco-mMiniLMv2-L12-H384-v1). Unlike bi-encoders (like E5), the Cross-Encoder processes the query and document simultaneously, capturing finer semantic nuances at a higher computational cost. This setup tests whether adding a deep semantic layer on top of the entity signal yields further improvements.
  4. 4.
    Standalone Cross-Encoder. To isolate the effect of the neural re-ranker and serve as a strict baseline for State-of-the-Art semantic retrieval without entity enhancement, we implemented a pure Cross-Encoder pipeline. The candidates retrieved directly by the dense vector index are re-ranked by the Cross-Encoder model, bypassing any entity-based signal or RRF fusion. The top candidates are then passed to the generator. This configuration allows us to verify whether the domain-specific entity signal adds distinct value compared to a generic, high-performance deep learning re-ranker.
@@
Figure 1 illustrates the complete workflow of our proposed ELERAG method. The architecture processes the query in parallel streams—extracting entity-based features and computing dense embeddings—before fusing the results via the RRF module to feed the LLM. The other experimental configurations can be understood as variations of this schema: the Standard RAG baseline utilizes only the lower branch (Dense Embedding $\rightarrow$ Vector Index $\rightarrow$ LLM), bypassing the Entity Linking and Re-ranking stages. The Weighted-Score Re-ranking configuration replaces the RRF block with the linear combination strategy described above. The RRF + Cross-Encoder Re-ranking configuration adds a further refinement block between the RRF stage and the LLM generation. Finally, the Standalone Cross-Encoder utilizes only the lower branch (Dense Embedding $\rightarrow$ Vector Index $\rightarrow$ Cross-Encoder $\rightarrow$ LLM) bypassing the Entity Linking and Re-ranking stages and adding the Cross-Encoder step.
@@ 3.1 Educational Data
The dataset used in this study was constructed from two Italian university courses, namely Applied Economics and Language and Communication, offered by an Italian telematic university. Italian was selected as the target language because it allowed the construction of a novel evaluation dataset derived from educational lecture material, which differs significantly from widely used public QA benchmarks. From a resource perspective, Italian also represents an intermediate scenario: while supported by Wikipedia and Wikidata, its coverage is substantially smaller than English, making it suitable for evaluating the robustness of the proposed approach outside extremely resource-rich settings. Each course consisted of several video lectures, totalling 50 classes over 32 hours of material, which were first transcribed into text using the Whisper Turbo automatic speech recognition model (radford2022whisper). The transcription process also generated precise word-level timestamps, which were later used to associate each text segment with its corresponding temporal interval in the original lecture.
To prepare the textual material for retrieval, the transcribed lectures were segmented into coherent chunks using the SpaCy Python library (spacy2023). To ensure semantic coherence, segmentation was strictly constrained to sentence boundaries: sentences were grouped to form chunks with a length between 20 and 300 tokens. This strategy prevents cutting sentences in the middle, preserving the syntactic and semantic integrity of the text. Crucially, because the chunking process respects sentence boundaries, the local context required for Named Entity Recognition remains fully intact. Consequently, when the stored chunks are subsequently processed by the SpaCy NER pipeline (prior to the Entity Linking module described in Section 2.2), the quality of entity extraction is not negatively affected by arbitrary word-level cuts.
To quantify the density of factual and domain-specific information within the corpus, we computed the entity coverage metrics following the offline Entity Linking phase. Out of a total of 676 chunks, 540 chunks contain at least one successfully linked Wikidata entity, yielding an overall chunk coverage of 79.88%. In total, 4,393 entities were extracted and disambiguated across the corpus, averaging 6.50 linked entities per chunk. This high entity density confirms that the educational narrative is rich in specific terminology, concepts, and factual references, providing a solid foundation for the subsequent entity-based re-ranking phase.
@@ 3.2 Evaluation setup
After generation, the dataset underwent a two-step validation process. First, a secondary GPT-4o prompt was used to verify the correctness and consistency between questions and their corresponding answers, automatically filtering invalid records. Subsequently, a manual curation was performed collaboratively by the authors, acting as domain experts, to identify and filter out any residual ambiguous, ill-formed, or hallucinated questions that might have bypassed the automated check. Because this final refinement was conducted as an expert consensus process rather than a parallel labeling task by independent annotators, standard inter-annotator agreement metrics were not applicable. All authors reviewed the generated questions and reached full consensus on the final selection. This rigorous two-step process ensured the high quality of the final set of 69 questions used for evaluation.
@@ 3.3 Evaluation Metrics
  • Subjective metrics: Completeness, relevance, and clarity, each evaluated by gpt-4o on a 1–10 scale for the generated answers obtained by concatenating the top-3 retrieved chunks. In this setup, the evaluator LLM is provided only with the retrieved chunks and does not have access to the generated answers. This design ensures that the evaluation focuses on the informational quality of the retrieved passages and avoids potential feedback loops or self-evaluation bias. Although LLM-based evaluation cannot fully replace human judgment, it provides a scalable and replicable proxy for consistently estimating answer quality across the benchmark.
@@ 3.4 Evaluation methods
    Method 2: Subjective evaluation through LLM-based scoring. To assess the informational quality of the retrieval, the top-$K$ retrieved chunks ($K=3$) were directly concatenated into a single text block. A separate gpt-4o instance then acted as an external evaluator, scoring these raw human transcripts on a 1–10 scale for completeness, relevance, and clarity. Crucially, because the evaluated text is not LLM-generated, this setup is inherently immune to stylistic “self-preference bias”. Because the evaluated text consists exclusively of concatenated retrieved passages rather than generated answers, the evaluator LLM does not assess its own outputs. This prevents feedback loops and mitigates the well-known self-preference bias observed in LLM-as-a-judge evaluations. The exact prompt is available on GitHub  [^5] https://github.com/Granataaa/educational-rag-el/blob/main/rag_api/testBench2.py.
@@ 4 Results
  • Hybrid Sparse-Dense Baseline (Dense + BM25): This hybrid approach is also applied to the Top-30 candidate chunks retrieved by FAISS. Note that this baseline was introduced specifically to isolate pure retrieval performance and was exclusively evaluated under Method 1 using our custom educational dataset.
@@
  • Cross-Encoder (Standalone): Re-scoring of Top-30 candidates retrieved by FAISS.
@@ 4.1 Method 1: Classical Evaluation on Retrieval
Cross-Encoder 0.536 0.408 0.633 0.693 0.816 0.652 0.646 0.755
@@
Our proposed ELERAG strategy emerged as the most effective method for this domain, achieving the highest scores in key precision metrics (EM, MRR, P@1). This superiority demonstrates that fusing the entity signal with dense retrieval via RRF is more robust than a simple linear combination (Weighted-Score), effectively filtering out semantically similar but factually incorrect chunks.
The Standalone Cross-Encoder achieved the highest Recall@10 ($0.816$), confirming the ability of deep transformer models to identify relevant content within a broader window. However, this broad recall did not translate into superior ranking precision: both Cross-Encoder variants scored lower than ELERAG in EM and MRR (both MRR_gold and MRR_rel_docs). This might depend on the fact that while the generic, pre-trained Cross-Encoder captures general semantic relevance effectively, it lacks the specific domain grounding provided by the explicit Entity Linking signal. In this specialized educational context, the Cross-Encoder tends to assign high scores to semantically plausible but factually distinct chunks, whereas ELERAG successfully utilizes Wikidata IDs to disambiguate and prioritize the exact target concept.
It is worth noting that the Hybrid Sparse-Dense Baseline and the Standalone Cross-Encoder maintain higher Recall@5 and Recall@10 scores, respectively, compared to ELERAG. This might be due to the fact that while entity linking can retrieve additional relevant passages, it may also introduce some noise, as passages discussing related entities or topics may be retrieved even when they are not directly relevant to the query. The RRF step mitigates this effect by down-weighting passages that are not consistently ranked highly across signals. However, when evaluating Recall@k with high values of k, some of these additional candidates may occupy relevant positions, which can slightly reduce recall compared to other baselines. This highlights a characteristic behavior of the RRF algorithm when applied to entity-based signals: it aggressively optimizes the very top positions (Top-1) to maximize precision for the limited context window of the LLM. Although this strict filtering may push some marginally relevant documents further down the list, the substantial gain in MRR and Exact Match is far more valuable for a RAG system, where the generator’s performance depends primarily on the quality of the very first retrieved chunks.
Furthermore, this comparison perfectly validates the architectural choice of integrating Entity Linking over traditional sparse retrieval. While the Hybrid Sparse-Dense Baseline (Dense+BM25) performs competitively in those mid-tier recall metrics, ELERAG significantly outperforms it in the most critical ranking metrics, specifically Precision@1 ($0.695$ vs. $0.637$) and both MRR metrics. This empirical evidence demonstrates that our Entity Linking module successfully goes beyond mere syntactic overlap: by grounding terms to Wikidata, it resolves disambiguations and captures domain-specific relationships that push the precise gold chunk to the very top of the ranking, which is an essential requirement for the subsequent generation phase.
To assess the dependence on the specific entity linking model, we also conducted experiments by varying the parameter $\alpha$ (see Sect. 2.2). Note that when $\alpha=0$, the ranking is based only on the Wikidata API score (i.e., popularity), while when $\alpha=1.0$, the ranking relies exclusively on the similarity score. The results are reported in Figure 2.
Figure 2: Ablation Study of $\alpha$ parameter in the Entity Disambiguation module. The plot shows that between 0.0 and 1.0 there are almost no variation of EM and MRR.
We observed almost no variation when $\alpha$ ranges between 0.0 and 1.0 (EM consistently equals 0.565, while MRR ranges between 0.668 and 0.674), confirming that our method is largely independent of the specific entity linking strategy employed.